during a high school spirit week student dress up in costumes. a cash price is given to students with the best costume the organization committee has $1000 to give away over 5 days the committee wants to increase the price by $50 each day how much should the committee give away on the first day?

1
Expert's answer
2022-06-02T17:15:31-0400

Since the committee wants to increase the price by $50 each day, the given money is the arithmetic progression with the common difference "d=50."


Since the high school spirit week lasts 5 days, we have 5 terms of arithmetic progression:

"a_1, a_2, ..., a_5" , and their sum: "S_5=1000."


"S_5=\frac{5(a_1+a_5)}{2}" , and "a_5=a_1+4d=a_1+4\cdot 50=a_1+200,"

so we have:

"1000=\frac{5(2a_1+200)}{2}=\frac{5\cdot 2(a_1+100)}{2}=5(a_1+100),"


"a_1+100=1000 : 5=200,"

"a_1=200-100,"

"a_1=100."


We have next prices over 5 days: $100, $150, $200, $250, $300.

And "100+150+200+250+300=1000."


Answer: the committee should give away $100 on the first day